# PanoDreamer: 3D Panorama Synthesis from a Single Image

### v1

@@ Abstract
In this paper, we present PanoDreamer, a novel method for producing a coherent 360° 3D scene from a single input image. Unlike existing methods that generate the scene sequentially, we frame the problem as single-image panorama and depth estimation. Once the coherent panoramic image and its corresponding depth are obtained, the scene can be reconstructed by inpainting the small occluded regions and projecting them into 3D space. Our key contribution is formulating single-image panorama and depth estimation as two optimization tasks and introducing alternating minimization strategies to effectively solve their objectives. We demonstrate that our approach outperforms existing techniques in single-image 360° scene reconstruction in terms of consistency and overall quality111people.engr.tamu.edu/nimak/Papers/PanoDreamer.
@@
Figure 1: We introduce a novel method for 360° 3D scene synthesis from a single image. Our approach generates a panorama and its corresponding depth in a coherent manner, addressing limitations in existing state-of-the-art methods such as LucidDreamer [3] and WonderJourney [36]. These methods sequentially add details by following a generation trajectory, often resulting in visible seams when looping back to the input image. In contrast, our approach ensures consistency throughout the entire 360° scene, as shown. The yellow bars show the regions corresponding to the input in each result.
1 Introduction
@@
Generating immersive and realistic 3D scenes from a single input image has emerged as one of the important topics in computer vision/graphics, driven by its broad applications including virtual/augmented reality (VR/AR) and gaming. While early algorithms [17, 22, 12, 10, 42, 24, 15, 25] have achieved high-quality results, they are generally limited to synthesizing novel views with only minor deviation from the input camera position. Consequently, these techniques cannot reconstruct a full 360° scene, which is the primary goal of our work.
@@
With the introduction of diffusion models, the more recent approaches have focused on utilizing these powerful models for 3D scene reconstruction. Specifically, several methods [16, 41, 40] propose various ways to generate 3D scenes from an input text prompts. These methods first generate entire panorama from text prompt using pretrained text-to-panorama diffusion models (DMs) and then lift it to 3D. Unfortunately, these approaches are fully generative and do not have a mechanism for reconstructing a 3D scene which is also consistent with a single input image.
@@
Several methods [3, 23, 36, 18, 8, 37] specifically address the problem of 3D scene reconstruction from a single image. Starting from the input image, these methods typically project it into 3D space, render it from a novel view, and then inpaint the missing regions using a diffusion model. They repeat this process for a series of cameras along a specific path to reconstruct the complete 3D scene. However, a major limitation of these approaches is that, due to the progressive nature of the scene building, they often fail to synthesize coherent 360° scenes, i.e., the start and end of the 360° scenes are contextually different.
@@
In this work, we propose a novel framework, coined PanoDreamer, for generating a coherent 360° 3D scene from a single input image. Departing from the existing methods, which generate the 3D scene one image at a time, we start by producing a single coherent 360° panorama from the input image using standard pre-trained inpainting diffusion models. Inspired by MultiDiffusion [1], we formulate the problem as an optimization with two loss terms and propose an alternating minimization strategy to optimize the objective, resulting in a coherent and seamless panoramic image.
@@
The next stage of our approach involves estimating the depth of the panoramic image to project pixels into 3D space and reconstruct the 3D scene. While powerful monocular depth estimation methods [31] exist, these techniques are typically optimized for specific resolutions and struggle to handle large panoramic images effectively. To address this problem, we formulate panoramic depth reconstruction as an optimization task, aiming to simultaneously produce a coherent panoramic depth map and a parametric function that aligns the range of monocular depth to the target depth. We propose an alternating minimization approach to efficiently solve this objective, resulting in a coherent and seamless panoramic depth map.
@@
Given the panoramic image and depth, we directly apply the approach of Shih et al. [22] to construct a layered depth image (LDI) and inpaint the missing regions in each layer. Next, we build a 3D Gaussian splatting (3DGS) representation[11] by initializing a set of Gaussians through the projection of LDI pixels into 3D space. We then optimize the 3DGS representation to sharpen details and obtain the final scene. We demonstrate that PanoDreamer can reconstruct consistent 360° 3D scenes from single input images that outperform existing methods. In summary, our work makes the following contributions:
@@
2 Related Work
2.1 Panorama Generation
@@
Diffusion models (DMs) have shown promising results across various generative tasks. In particular, several approaches [1, 13, 4, 33, 14, 6, 26, 38] have proposed leveraging pretrained DMs to synthesize panoramic images. For example, DiffCollage [38] reconstructs complex factor graphs and aggregates intermediate output from DMs defined by nodes to generate a panorama. PanoGen [14] utilizes latent diffusion models combined with recursive outpainting to create indoor panoramic images. MultiDiffusion [1] frames the problem of panoramic image generation from pretrained DMs as an optimization process to produce globally consistent images. SynDiffusion [13] builds on this idea and incorporates the LPIPS score [39] between neighboring denoised images into the optimization process. StitchDiffusion [26] further proposes averaging the overlapping denoising predictions and fine-tuning a low-rank adaptation (LoRA) module [9]. To improve the efficiency of the generation process, SpotDiffusion [6] shifts non-overlapping denoising windows over time to synthesize a coherent panorama in an efficient manner. All of these methods generate panoramas purely from a text prompt and cannot incorporate an input image into the generation process.
@@
In contrast to these approaches, PanoDiffusion [30] is designed to generate panoramas from a masked input image. Similarly, MVDiffusion [4] can produce a panorama from a single image by stitching multiple images using pixel-wise correspondences and attention modules. However, both of these approaches require training and struggle to generalize to diverse scenarios.
2.2 View synthesis from a single input image
@@
Numerous methods have been proposed to synthesize scenes from a single input image. One category of these methods [17, 22, 12, 10] addresses this problem in a modular manner, decomposing the synthesis process into several independent components. For example, Shih et al.[22] estimate a layered depth image (LDI) representation to reconstruct novel views. Niklaus et al.[17] use the estimated depth to map the input image to a point cloud and train a network to fill in the missing areas.
@@
The second group of methods [42, 24, 15, 25] synthesizes scenes from a single input image in an end-to-end manner. Among these approaches, Zhou et al.[42] propose synthesizing scenes by first estimating optical flow and then warping the input image to novel views. Srinivasan et al.[24] use two sequential convolutional neural networks to estimate disparity and refine the warped images. Several approaches propose synthesizing intermediate scene representations to achieve view synthesis. For example, SynSin [28] estimates a point cloud of a scene, and several methods [15, 25] synthesize light fields using the estimated Multiple Plane Image (MPI) representation. PixelNeRF [34] trains a NeRF prior and can synthesize NeRF from a single input image without performing test-time optimization. Additionally, several approaches [19, 2] focus on improving the view-dependent effects for single-view view synthesis. However, all of these methods are designed only for view synthesis within a narrow viewing angle or restricted camera movement and cannot be generalized well to the entire 360° scene.
2.3 3D Scene Generation
@@
The first group of methods [18, 23, 37, 5, 36, 35, 3] generates 3D scenes from text or images in a progressive manner. Starting from a single image, either provided by the user or generated from a text prompt, these methods typically perform progressive inpainting, monocular depth estimation, and 3D optimization for novel views in the 3D scene. These approaches differ in their 3D representation, image inpainting, and depth refinement strategies. However, since the 3D scene is generated through progressive inpainting of single inputs, these methods struggle to preserve coherency, making it difficult to synthesize consistent 360° scenes.
The second group of methods [16, 41, 40] generates 360° 3D scenes in a two-step process. They synthesize coherent panoramas by leveraging pretrained text-to-panorama DMs, which are then lifted to 3D using different inpainting and depth estimation strategies. Although these approaches are capable of generating consistent 3D scenes from single inputs, they are text-conditioned only and do not have any mechanism to reconstruct a scene consistent with a single input image. In comparison, our method, PanoDreamer, not only generates coherent 3D scenes but also allows users to condition the generation on any single input image.
3 Preliminaries
In this section, we describe MultiDiffusion [1], an approach that leverages a pre-trained diffusion model, without any fine-tuning, to produce results in various image or condition spaces. For example, this technique can generate outputs at resolutions different from the base model’s native resolution (e.g., panoramas) or synthesize images using region-based text prompts. Here, we focus our discussion on the former example, as it is most relevant to our approach.
@@ 3 Preliminaries
$I_{t-1}=\Phi(I_{t}|p).$ (1)
The goal of MultiDiffusion is to leverage this base model to generate an image at a larger resolution $H^{\prime}\times W^{\prime}$. The MultiDiffusion process, similarly, begins with a noisy high-resolution image, $J_{T}$, and produces a clean image $J_{0}$ through a sequence of gradually denoised images $J_{T-1},\cdots,J_{0}$. At each time step, given the optimal high-resolution image $J^{*}_{t}$, the key idea is to ensure that the next step, $J_{t-1}$, aligns closely with the output of the base diffusion model. Specifically, the desired $J_{t-1}$ ensures that $F_{i}(J_{t-1})\sim\Phi(F_{i}(J^{*}_{t})|p)$, where $F_{i}$ is an operator that maps the high-resolution image space to the base model’s space (via cropping, in this case). Enforcing this similarity in the $L_{2}$ sense, we arrive at the following objective:
$J^{*}_{t-1}=\arg\min_{J}\sum_{i=1}^{n}\left\|W_{i}\odot\left[F_{i}(J)-\Phi(F_{i}(J^{*}_{t})|p)\right]\right\|^{2},$ (2)
where $W_{i}$ is a weight map ($W_{i}=\mathbf{1}$ in this case) and $\odot$ denotes the element-wise product.
@@
$J^{*}_{t-1}=\sum_{i=1}^{n}\frac{F_{i}^{-1}(W_{i})}{\sum_{i=1}^{n}F_{i}^{-1}(W_{i})}\odot F_{i}^{-1}(\Phi(F_{i}(J^{*}_{t})|p)),$ (3)
@@
where $F_{i}^{-1}$ is the inverse of the cropping operator, which places the content into appropriate location in the high-resolution image. Starting from the noisy high-resolution image $J^{*}_{T}=J_{T}$, MultiDiffusion uses this process to obtain the optimal intermediate high-resolution images $J^{*}_{t}$, resulting in the final clean image $J^{*}_{0}$.
4 Algorithm
@@
Given a single input image $I$, our goal is to reconstruct a coherent 360 scene using a 3D Gaussian representation [11]. Unlike existing methods that produce the 3D scenes through progressive projection and inpainting, we begin by generating a coherent 360° panorama from the input image (Sec. 4.1). We then estimate a coherent and consistent depth from the generated panorama (Sec. 4.2). Finally, we inpaint the occluded regions using layered depth image (LDI) inpainting and use the inpainted layers to reconstruct a 3DGS representation (Sec. 4.3).
Figure 2: We compare the results of our MultiConDiffusion process against MultiDiffusion and progressive inpainting. The green bar shows the location of the input image.
4.1 Single-Image Panorama Generation
@@
We begin by discussing the problem of generating a larger image from a single input image, then explain the specific details for panorama generation. Given an input image $I$ placed on a larger canvas $L$ of size $H^{\prime}\times W^{\prime}$, our goal is to fill in missing areas in $L$ using an inpainting diffusion model $\Phi$, which operates on fixed lower-resolution images of size $H\times W$. In addition to a text prompt $p$, this model takes a mask $M$ denoting the missing regions and a masked image $(1-M)\odot X$ as inputs. It progressively denoises a Gaussian noise image $I_{T}$ to obtain a clean image $I_{0}$ containing the hallucinated details, with each step following $I_{t-1}=\Phi(I_{t}|p,M,(1-M)\odot X)$. A straightforward approach is to use this model to gradually outpaint the high-resolution image, starting from the regions covered by the input. However, this approach often results in noticeable contextual inconsistencies and seams, as shown in Fig. 2.
Inspired by MultiDiffusion, we address this issue by formulating the problem as an optimization task. Similar to MultiDiffusion, we aim to optimize a sequence of iteratively denoised high-resolution images $J_{T-1},\dots,J_{0}$, beginning with the noisy image $J_{T}$. To construct our objective, we first discuss our two desired properties. First, given the optimal denoised high-resolution image at the current time $J^{*}_{t}$, the image at the next time $J_{t-1}$ should closely approximate the output of the base inpainting model, i.e., $F_{i}(J_{t-1})\sim\Phi(F_{i}(J^{*}_{t})|p,M_{ij},M_{ij}\odot F_{i}(L))$. This property is analogous to that in MultiDiffusion, but with the distinction that here, the similarity holds not only for random crops defined by $F_{i}$ but also for random inpainting masks within each crop, $M_{ij}$.
The first property alone is insufficient for producing high-quality results, as shown in Fig. 2 (MultiDiffusion). Unlike MultiDiffusion, our approach conditions the process on a masked image, $M_{ij}\odot F_{i}(L)$. Starting with a fixed initial $L$ (missing content everywhere except in the input), the inpainting model must hallucinate entire patches when $F_{i}$ does not overlap with the input region, leading to disconnection from the original input. To address this, we treat $L$ as a parameter to be optimized during our process. To set up the objective, we consider what would be ideal for $L$ in this context. If the clean high-resolution image $J_{0}$ were available, the best possible condition for guiding the diffusion process to reach $J_{0}$ would be to set $L$ to $J_{0}$. Armed by this observation, we establish our second property: ensuring that $L\sim J_{0}$.
Figure 3: We show the results of MultiConDiffusion during different iterations of the optimization.
By formalizing these two properties in the $L_{2}$ sense, we arrive at the following objective:
$\tilde{J}_{0}\mathinner{\cdotp\mkern-2.0mu\cdotp\mkern-2.0mu\cdotp}\tilde{J}_{T-1},\tilde{L}=\operatorname*{arg\,min}_{J_{0}\mathinner{\cdotp\mkern-2.0mu\cdotp\mkern-2.0mu\cdotp}J_{T-1},L}\left[\sum_{t=T}^{1}\mathcal{L}(J_{t-1}|J^{*}_{t})+\|L-J_{0}\|^{2}\right]$ (4)
@@
$\mathcal{L}(J|J^{*}_{t})=\sum_{j=1}^{m}\sum_{i=1}^{n}\|M_{ij}\odot\left[F_{i}(J_{t-1})-\Phi(F_{i}(J^{*}_{t})|\mathcal{C}_{ij})\right]\|^{2},$ (5)
and
$\mathcal{C}_{ij}=\{p,M_{ij},M_{ij}\odot F_{i}(L)\}.$ (6)
Here, the first term ensures that the cropped image at the next time step, $F_{i}(J_{t-1})$, is similar to the output of the base inpainting diffusion model in the masked regions $M_{ij}$. The second term simply forces the condition image $L$ to be close to the clean high-resolution image $J_{0}$. Note that our high-resolution diffusion process, called MultiConDiffusion, is also conditional. The output of this process, $J_{T-1},\dots,J_{0}$, depends on the condition image $L$. As such, $J^{*}_{t}$ is the optimal solution at time $t$ given the current condition image $L$, and it is different from the final optimal solution, $\tilde{J}_{t}$, which is obtained given the optimal condition image $\tilde{L}$.
@@
Here, we fix $L$ and minimize Eq.4 by finding the optimal $J_{T-1},\dots,J_{0}$. Since $J_{T-1},\dots,J_{1}$ do not influence the second term (different steps are assumed to be independent), we can simply use Eq.3 to obtain their solution in closed form. The main difference here is that the base diffusion model is replaced with the inpainting model, and $W_{i}=M_{ij}$. On the other hand, since $J_{0}$ appears in both terms and they are both quadratic with respect to it, the final solution is a weighted combination of the solution to the first term (Eq. 3) and the second term ($J^{*}_{0}=L$). In practice, however, we found that plausible results can still be obtained even when ignoring the second term.

### v3

# PanoDreamer: Optimization-Based Single Image to 360 3D Scene With Diffusion

@@ Abstract.
In this paper, we present PanoDreamer, a novel method for producing a coherent 360° 3D scene from a single input image. Unlike existing methods that generate the scene sequentially, we frame the problem as single-image panorama and depth estimation. Once the coherent panoramic image and its corresponding depth are obtained, the scene can be reconstructed by inpainting the small occluded regions and projecting them into 3D space. Our key contribution is formulating single-image panorama and depth estimation as two optimization tasks and introducing alternating minimization strategies to effectively solve their objectives. We demonstrate that our approach outperforms existing techniques in single-image 360° 3D scene reconstruction in terms of consistency and overall quality111people.engr.tamu.edu/nimak/Papers/PanoDreamer.
Single Image to 3D, 3D Scene Generation, Diffusion Models, Panorama Generation, Panorama Depth
submissionid: 1483journalyear: 2025copyright: ccconference: SIGGRAPH Asia 2025 Conference Papers; December 15–18, 2025; Hong Kong, Hong Kongbooktitle: SIGGRAPH Asia 2025 Conference Papers (SA Conference Papers ’25), December 15–18, 2025, Hong Kong, Hong Kongdoi: 10.1145/3757377.3763883isbn: 979-8-4007-2137-3/2025/12ccs: Computing methodologies Image-based rendering
@@
Figure 1. We introduce a novel method for 360° 3D scene synthesis from a single image. Our approach generates a panorama and its corresponding depth in a coherent manner, addressing limitations in existing state-of-the-art methods such as LucidDreamer (Chung et al., 2023) and WonderJourney (Yu et al., 2024b). These methods sequentially add details by following a generation trajectory, often resulting in visible seams when looping back to the input image. In contrast, our approach ensures consistency throughout the entire 360° scene. The yellow bars show the regions corresponding to the input in each result.
1. Introduction
@@
Generating immersive and realistic 3D scenes from a single input image has emerged as one of the important topics in computer vision/graphics, driven by its broad applications including virtual/augmented reality (VR/AR) and gaming. While early algorithms (Niklaus et al., 2019; Shih et al., 2020; Kopf et al., 2020; Jampani et al., 2021; Zhou et al., 2016; Srinivasan et al., 2017; Li and Kalantari, 2020; Tucker and Snavely, 2020) have achieved high-quality results, they are generally limited to synthesizing novel views with only minor deviation from the input camera position. Consequently, these techniques cannot reconstruct a full 360° scene, which is the primary goal of our work.
@@
With the introduction of diffusion models, the more recent approaches have focused on utilizing these powerful models for 3D scene reconstruction. Specifically, several methods (Li et al., 2024; Zhou et al., 2025, 2024) propose various ways to generate 3D scenes from input text prompts. These methods first generate entire panorama from text prompt using pretrained text-to-panorama diffusion models (DMs) and then lift it to 3D. Unfortunately, these approaches are fully generative and do not have a mechanism for reconstructing a 3D scene which is also consistent with a single input image.
@@
Several methods (Chung et al., 2023; Shriram et al., 2024; Yu et al., 2024b; Ouyang et al., 2023; Höllein et al., 2023; Zhang et al., 2024) specifically address the problem of 3D scene reconstruction from a single image. Starting from the input image, these methods typically project it into 3D space, render it from a novel view, and then inpaint the missing regions using a diffusion model. They repeat this process for a series of cameras along a specific path to reconstruct the complete 3D scene. However, a major limitation of these approaches is that, due to the progressive nature of the scene building, they often fail to synthesize coherent 360° scenes, i.e., the start and end of the 360° scenes are contextually different.
@@
In this work, we propose a novel framework, coined PanoDreamer, for generating a coherent 360° 3D scene from a single input image. Departing from the existing methods, which generate the 3D scene one image at a time, we start by producing a coherent 360° panorama from the input image using standard pre-trained inpainting diffusion models. Inspired by MultiDiffusion (Bar-Tal et al., 2023), we formulate the problem as an optimization with two loss terms and propose an alternating minimization strategy to optimize the objective, resulting in a coherent and seamless panoramic image.
@@
The next stage of our approach involves estimating the depth of the panoramic image to project pixels into 3D space and reconstruct the 3D scene. While powerful monocular depth estimation methods (Yang et al., 2024a) exist, these techniques are typically optimized for specific resolutions and struggle to handle large panoramic images effectively. To address this problem, we formulate panoramic depth reconstruction as an optimization task, aiming to simultaneously produce a coherent panoramic depth map and a parametric function that aligns the range of monocular depth to the target depth. We propose an alternating minimization approach to efficiently solve this objective, resulting in a coherent and seamless panoramic depth map.
@@
Given the panoramic image and depth, we directly apply the approach of Shih et al. (2020) to construct a layered depth image (LDI) and inpaint the missing regions in each layer. Next, we build a 3D Gaussian splatting (3DGS) representation (Kerbl et al., 2023) by initializing a set of Gaussians through the projection of LDI pixels into 3D space. We then optimize the 3DGS representation to sharpen details and obtain the final scene. We demonstrate that PanoDreamer can reconstruct consistent 360° 3D scenes from single input images that outperform existing methods. In summary, our work makes the following contributions:
@@
2. Related Work
2.1. Panorama Generation
@@
Diffusion models (DMs) have shown promising results across various generative tasks. In particular, several approaches (Bar-Tal et al., 2023; Lee et al., 2023; Deng et al., 2023; Ye et al., 2024; Li and Bansal, 2023; Frolov et al., 2024; Wang et al., 2024; Zhang et al., 2023) have proposed leveraging pretrained DMs to synthesize panoramic images. For example, DiffCollage (Zhang et al., 2023) reconstructs complex factor graphs and aggregates intermediate output from DMs defined by nodes to generate a panorama. PanoGen (Li and Bansal, 2023) utilizes latent diffusion models combined with recursive outpainting to create indoor panoramic images. MultiDiffusion (Bar-Tal et al., 2023) frames the problem of panoramic image generation from pretrained DMs as an optimization process to produce globally consistent images. SynDiffusion (Lee et al., 2023) builds on this idea and incorporates the LPIPS score (Zhang et al., 2018) between neighboring denoised images into the optimization process. StitchDiffusion (Wang et al., 2024) further proposes averaging the overlapping denoising predictions and fine-tuning a low-rank adaptation (LoRA) module (Hu et al., 2021). To improve the efficiency of the generation process, SpotDiffusion (Frolov et al., 2024) shifts non-overlapping denoising windows over time to synthesize a coherent panorama efficiently. All of these methods generate panoramas from a text prompt and cannot incorporate an input image into the generation process.
@@
In contrast to these approaches, PanoDiffusion (Wu et al., 2023b) is designed to generate panoramas from a masked input image. Similarly, MVDiffusion (Deng et al., 2023) can produce a panorama from a single image by stitching multiple images using pixel-wise correspondences and attention modules. However, both of these approaches require training and struggle to generalize to diverse scenarios.
2.2. View synthesis from a single input image
@@
Numerous methods have been proposed to synthesize scenes from a single input image. One category of these methods (Niklaus et al., 2019; Shih et al., 2020; Kopf et al., 2020; Jampani et al., 2021; Pu et al., 2023) addresses this problem in a modular manner, decomposing the synthesis process into several independent components. For example, Shih et al. (2020) estimate a layered depth image (LDI) representation to reconstruct novel views. Niklaus et al.(2019) use the estimated depth to map the input image to a point cloud and train a network to fill in the missing areas.
@@
The second group of methods (Zhou et al., 2016; Srinivasan et al., 2017; Li and Kalantari, 2020; Tucker and Snavely, 2020) synthesizes scenes from a single input image in an end-to-end manner. Among these approaches, Zhou et al.(2016) propose synthesizing scenes by first estimating optical flow and then warping the input image to novel views. Srinivasan et al.(2017) use two sequential convolutional neural networks to estimate disparity and refine the warped images. Several approaches propose synthesizing intermediate scene representations to achieve view synthesis. For example, SynSin (Wiles et al., 2020) estimates a point cloud of a scene, and several methods (Li and Kalantari, 2020; Tucker and Snavely, 2020) synthesize light fields using the estimated multi-plane image (MPI) representation. PixelNeRF (Yu et al., 2021) trains a NeRF prior and can synthesize NeRF from a single input image without performing test-time optimization. Additionally, several approaches (Paliwal et al., 2023; Bello and Kim, 2024) focus on improving the view-dependent effects for single-view view synthesis. However, all of these methods are designed only for view synthesis within a narrow angle or restricted camera movement and cannot be generalized to the entire 360° scene.
2.3. 3D Scene Generation
@@
The first group of methods (Ouyang et al., 2023; Shriram et al., 2024; Zhang et al., 2024; Engstler et al., 2024; Yu et al., 2024b, a; Chung et al., 2023; Liang et al., 2025; Ni et al., 2025) generates 3D scenes from text or images in a progressive manner. Starting from a single image, either provided by the user or generated from a text prompt, these methods typically perform progressive inpainting, monocular depth estimation, and 3D optimization for novel views in the 3D scene. These approaches differ in their 3D representation, image inpainting, and depth refinement strategies. However, since the 3D scene is generated through progressive inpainting of single inputs, these methods struggle to preserve coherency, making it difficult to synthesize consistent 360° scenes.
The second group of methods (Li et al., 2024; Zhou et al., 2025, 2024) generates 360° 3D scenes in a two-step process. They synthesize coherent panoramas by leveraging pretrained text-to-panorama DMs, which are then lifted to 3D using different inpainting and depth estimation strategies. Although these approaches are capable of generating consistent 3D scenes from inputs, they are text-conditioned only and do not have any mechanism to reconstruct a scene consistent with a single input image. In comparison, our method, PanoDreamer, not only generates coherent 3D scenes but also allows users to condition the generation on any single input image.
3. Preliminaries
In this section, we describe MultiDiffusion (Bar-Tal et al., 2023), an approach that leverages a pre-trained diffusion model, without any fine-tuning, to produce results in various image or condition spaces. For example, this technique can generate outputs at resolutions different from the base model’s native resolution (e.g., panoramas) or synthesize images using region-based text prompts. Here, we focus our discussion on the former example, as it is most relevant to our approach.
@@ 3. Preliminaries
(1) $I_{t-1}=\Phi(I_{t}|p).$
The goal of MultiDiffusion is to leverage this base model to generate an image $J_{0}$ at a larger resolution $H^{\prime}\times W^{\prime}$. The MultiDiffusion process begins with a noisy high-resolution image, $J_{T}$, and produces a clean image $J_{0}$ through a sequence of gradually denoised images $J_{T-1},\cdots,J_{0}$. Given the optimal high-resolution image at the current step, $J^{*}_{t}$, the key idea of MultiDiffusion is to ensure that the output of the base diffusion model $\Phi(F_{i}(J^{*}_{t})|p)$ is as close as possible to the high-resolution image at the next step $F_{i}(J_{t-1})$, locally. Note that $F_{i}$ is an operator that maps the high-resolution image space to the base model’s space (via cropping, in this case). Enforcing this similarity in the $L_{2}$ sense, we arrive at the following objective:
(2) $J^{*}_{t-1}=\arg\min_{J}\sum_{i=1}^{n}\left\|W_{i}\odot\left[F_{i}(J)-\Phi(F_{i}(J^{*}_{t})|p)\right]\right\|^{2},$
where $W_{i}$ is a weight map ($W_{i}=\mathbf{1}$ in this case), $n$ refers to the total number of crops, and $\odot$ denotes the element-wise product.
@@
(3) $J^{*}_{t-1}=\sum_{i=1}^{n}\frac{F_{i}^{-1}(W_{i})}{\sum_{j=1}^{n}F_{j}^{-1}(W_{j})}\odot F_{i}^{-1}(\Phi(F_{i}(J^{*}_{t})|p)),$
where $F_{i}^{-1}$ is the inverse of the cropping operator, which places the content into the appropriate location in the high-resolution image. At a high level, this solution aggregates (adds) the outputs of the base diffusion model for overlapping crops and normalizes the resulting image by the total number of crops at each pixel.
@@
Starting from the noisy high-resolution image $J^{*}_{T}=J_{T}$, MultiDiffusion uses this process to obtain the optimal intermediate high-resolution images $J^{*}_{t}$, resulting in the final clean high-resolution image $J^{*}_{0}$.
4. Algorithm
@@
Given a single input image $I$, our goal is to reconstruct a coherent 360 scene using a 3D Gaussian representation (Kerbl et al., 2023). Unlike existing methods that produce the 3D scenes through progressive projection and inpainting, we begin by generating a coherent 360° panorama from the input image (Sec. 4.1). We then estimate a coherent and consistent depth from the generated panorama (Sec. 4.2). Finally, we inpaint the occluded regions using layered depth image (LDI) inpainting and use the inpainted layers to reconstruct a 3DGS representation (Sec. 4.3).
4.1. Single-Image Panorama Generation
@@
We begin by discussing our method for generating a larger image from a single input image, then explain the specific details for panorama generation in Sec. 4.1.1. The overview of our approach is provided in Fig. 2. Given an input image $I$ placed on a larger canvas $L$ of size $H^{\prime}\times W^{\prime}$, our goal is to fill in missing areas in $L$ using an inpainting diffusion model $\Phi$, which operates on fixed lower-resolution images of size $H\times W$. In addition to a text prompt $p$, this model takes a mask $M$ denoting the missing regions and a masked image $(1-M)\odot X$ as inputs. It progressively denoises a Gaussian noise image $I_{T}$ to obtain a clean image $I_{0}$ containing the hallucinated details, with each step following $I_{t-1}=\Phi(I_{t}|p,M,(1-M)\odot X)$. A straightforward approach is to use this model to gradually outpaint the high-resolution image, starting from the regions covered by the input. However, this approach often results in noticeable contextual inconsistencies and seams, as shown in Fig. 3 (Progressive Inpainting).
Inspired by MultiDiffusion, we address this issue by formulating the problem as an optimization. MultiDiffusion can be adapted to this problem in a straightforward manner by replacing the base diffusion model with an inpainting model and reformulating the objective in Eq. 2 as follows:
(4) $\mathcal{L}(J_{t-1}|J^{*}_{t})=\hskip-3.61371pt\sum_{i=1}^{n}\|M_{i}\odot\left[F_{i}(J_{t-1})-\Phi(F_{i}(J^{*}_{t})|\mathcal{C}_{i})\right]\|^{2}\hskip-3.61371pt,$
@@
(5) $\mathcal{C}_{i}=\{p,M_{i},M_{i}\odot F_{i}(L)\}.$
Here, $M_{i}$ is a random inpainting mask for the $i^{\text{th}}$ crop, and $L$ is the high-resolution condition image. This objective ensures that the output of the inpainting diffusion model, $\Phi(F_{i}(J^{*}_{t})|\mathcal{C}_{i})$, is as close as possible to the corresponding crop of the high-resolution image at the next step $F_{i}(J_{t-1})$ in the masked areas $M_{i}$. This objective can be minimized similarly to Eq. 3 as follows:
(6) $J^{*}_{t-1}=\sum_{i=1}^{n}\frac{F_{i}^{-1}(M_{i})}{\sum_{j=1}^{n}F_{j}^{-1}(M_{j})}\odot F_{i}^{-1}(\Phi(F_{i}(J^{*}_{t})|\mathcal{C}_{i})),$
Figure 2. We provide an overview of our proposed MultiConDiffusion process, which consists of two stages. In the first stage, we fix the input condition $L$ and apply the diffusion model to overlapping crops of the image at the current time step. The outputs are then aggregated to produce the image at the next time step. This process is repeated until the fully denoised image $J_{0}$ is obtained. In the next stage, we replace the current input condition with $J_{0}$. These two stages are repeated until convergence.
Based on this equation, it is easy to infer that the solution depends on the high-resolution input condition, $L$, of the MultiDiffusion process. As shown in Fig. 3, MultiDiffusion results vary drastically depending on how the input condition is set. In particular, both simple methods for obtaining the input condition, such as placing the input image on a black canvas or using progressive inpainting, produce inconsistent results.
To address this issue, we make a key observation that the ideal input condition is a coherent and consistent high-resolution image. However, obtaining such an image, $J_{0}$, is the goal of our optimization and is not available beforehand. Therefore, we propose to incorporate this observation as an additional term in our objective as follows:
(7) $\tilde{J}_{0}\mathinner{\cdotp\mkern-2.0mu\cdotp\mkern-2.0mu\cdotp}\tilde{J}_{T-1},\tilde{L}=\operatorname*{arg\,min}_{J_{0}\mathinner{\cdotp\mkern-2.0mu\cdotp\mkern-2.0mu\cdotp}J_{T-1},L}\left[\sum_{t=T}^{1}\mathcal{L}(J_{t-1}|J^{*}_{t})+\|L-J_{0}\|^{2}\right]$}
where the first term is the adapted MultiDiffusion objective for all time steps, while the second term forces the condition image $L$ to be close to the clean high-resolution image $J_{0}$. Note that the output of this process, $J_{T-1},\dots,J_{0}$, depends on the condition image $L$. As such, $J^{*}_{t}$ is the optimal solution at time $t$ given the current condition image $L$, and it differs from the final optimal solution, $\tilde{J}_{t}$, which is obtained using the optimal condition image $\tilde{L}$. We call this equation the MultiConDiffusion objective, as the high-resolution diffusion process in our case is conditional.
Figure 3. We compare the results of our MultiConDiffusion process against MultiDiffusion and progressive inpainting. The green bar shows the location of the input image. We show MultiDifussion results with two different input conditions (shown on the top left): black canvas with input image (second row), and progressive inpainting result (third row). Our method produces coherent results, while the alternative approaches produce images with seams and inconsistencies.
@@
Here, we fix $L$ and minimize Eq. 7 by finding the optimal $J_{T-1},\dots,J_{0}$. Since $J_{T-1},\dots,J_{1}$ do not influence the second term (as different steps are assumed to be independent), we can use Eq. 6 to obtain their solution in closed form. On the other hand, since $J_{0}$ appears in both terms, and both terms are quadratic with respect to it, the final solution is a weighted combination of the solution to the first term (Eq. 6) and the second term ($J^{*}_{0}=L$). In practice, however, we found that plausible results can still be obtained even when the second term is ignored.
To summarize, as shown in Fig. 2, starting from $J_{T}$, we aggregate the output of the inpainting diffusion model over different overlapping crops to obtain the image at the next time step, resulting in a sequence of optimal $J^{*}_{T-1},\dots,J^{*}_{0}$ given the current fixed high-resolution input condition $L$.